4) A projectile is launched from the edge of a cliff with a velocity of 10.0 m/s at an angle of 30.0 degrees above the horizontal. What is the velocity of the projectile 2.00 seconds later (ignoring air resistance, and assuming it hasn't reached the ground yet)? A) VX-8.7 m/s, Vy -14.6 m/s B) VX = 14.6 m/s, Vy --5.6 m/s C) Vx = -8.7 m/s, Vy --14.6 m/s D) Vx--14.6 m/s, Vy - 8.7 m/s E) Vx = 5.6 m/s, Vy --14.6 m/s
1
Expert's answer
2020-02-12T10:40:48-0500
"V_x=V_0 \cos{30}"

"V_x=10 \cos{30}=8.7\frac{m}{s}"

"V_y=V_0 \sin{30}-gt"

"V_y=10 \sin{30}-(9.8)(2)=-14.6\frac{m}{s}"

Answer: A).